# Results related to the Gaussian product inequality conjecture for mixed-sign exponents in arbitrary dimension

### v1

@@ Abstract
This note establishes that the opposite Gaussian product inequality (GPI) of the type proved by Russell & Sun, (2022a) in two dimensions, and partially extended to higher dimensions by Zhou et al. , (2024), continues to hold for an arbitrary mix of positive and negative exponents. A general quantitative lower bound is also obtained conditionally on the GPI conjecture being true.
@@
centered Gaussian random vector , componentwise convex order , Gaussian product inequality , Loewner order , multivariate normal distribution , opposite Gaussian product inequality , real linear polarization constant
@@ 1 Introduction
The present note investigates the general mixed-sign problem, thereby significantly generalizing (3) and (4). Except for a narrow unresolved case discussed in Remark 2, the analysis resolves the question through Theorems 1 and 2.
@@ 2 Results
Next, the goal is to generalize (4) by replacing $|X_{d}|^{2\nu_{d}}$ with a product of positive powers of the absolute Gaussian components indexed in $\mathcal{J}^{\complement}$ or, more generally, a componentwise convex function of such components.
@@
Let $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma)$ for some positive definite matrix $\Sigma$, and let $\emptyset\neq\mathcal{J}\subseteq\{1,\ldots,d\}$ be given. Assume that $\nu_{j}\in[0,1/2)$ for all $j\in\mathcal{J}$. Then, for any componentwise convex function $\psi:\mathbb{R}^{d-|\mathcal{J}|}\to\mathbb{R}$, we have
@@
Since $(x_{i})_{i\in\mathcal{J}^{\complement}}\mapsto\prod_{i\in\mathcal{J}^{\complement}}|x_{i}|^{2\nu_{i}}$ is componentwise convex for all $\nu_{i}\in[1/2,\infty),~{}i\in\mathcal{J}^{\complement}$, the following corollary is immediate.
@@
Let $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma)$ for some positive definite matrix $\Sigma$, and let $\emptyset\neq\mathcal{J}\subseteq\{1,\ldots,d\}$ be given. Assume that $\nu_{j}\in[0,1/2)$ for all $j\in\mathcal{J}$ and $\nu_{i}\in[1/2,\infty)$ for all $i\in\mathcal{J}^{\complement}$. Then, we have
$\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\times\prod_{i\in\mathcal{J}^{\complement}}|X_{i}|^{2\nu_{i}}\right]\leq\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\right]\times\mathsf{E}\left[\prod_{i\in\mathcal{J}^{\complement}}|X_{i}|^{2\nu_{i}}\right].$
@@
The case where some of the exponents $\nu_{i},~{}i\in\mathcal{J}^{\complement},$ lie in the interval $(0,1/2)$ remains open. One possible approach would be to exploit the Bernstein function integral representation, $|x|^{2\nu}=\nu\int_{0}^{\infty}(1-e^{-sx^{2}})s^{-\nu-1}\mathrm{d}s/\Gamma(1-\nu)$, which holds for $x\in\mathbb{R}$ and $\nu\in(0,1/2)$; see, e.g., Schilling et al. , (2012, Eq. (1)). However, we were unable to turn this representation into a successful argument. A treatment of this case therefore lies beyond the scope of our paper.
@@ 3 Proofs
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]=\frac{1}{\prod_{j=1}^{k}\Gamma(\nu_{j})}\int_{(0,\infty)^{k}}\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]\prod_{j=1}^{k}s_{j}^{\nu_{j}-1}\mathrm{d}\boldsymbol{s},$ (6)
@@
$\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]$ $=\int_{(0,\infty)^{d}}\exp(-\boldsymbol{x}^{\top}T_{\boldsymbol{s}}\boldsymbol{x}/2)\,\frac{\exp(-\boldsymbol{x}^{\top}\Sigma^{-1}\boldsymbol{x}/2)}{(2\pi)^{d/2}|\Sigma|^{1/2}}\prod_{i=k+1}^{d}|x_{i}|^{2\nu_{i}}\mathrm{d}\boldsymbol{x}$
$=\frac{|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}{|\Sigma|^{1/2}}\int_{(0,\infty)^{d}}\frac{\exp\{-\boldsymbol{x}^{\top}(\Sigma^{-1}+T_{\boldsymbol{s}})\,\boldsymbol{x}/2\}}{(2\pi)^{d/2}|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}\prod_{i=k+1}^{d}|x_{i}|^{2\nu_{i}}\mathrm{d}\boldsymbol{x}$
$=\frac{|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}{|\Sigma|^{1/2}}\,\mathsf{E}\left[\prod_{i=k+1}^{d}|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right],$
@@
$\mathsf{Var}(Y^{(\boldsymbol{s})}_{i})=\boldsymbol{e}_{i-k}^{\top}\{(\Sigma/A)+B^{\top}A^{-1}(A^{-1}+S)^{-1}A^{-1}B\}\boldsymbol{e}_{i-k}\geq\boldsymbol{e}_{i-k}^{\top}(\Sigma/A)\boldsymbol{e}_{i-k}=\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\mathsf{Var}(X_{i}),$
@@
$\mathsf{E}\left[|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right]\geq\left\{\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\right\}^{\nu_{i}}\frac{\Gamma(\nu_{i}+1/2)}{\sqrt{\pi}}2^{\nu_{i}}\{\mathsf{Var}(X_{i})\}^{\nu_{i}}\\ =\left\{\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\right\}^{\nu_{i}}\mathsf{E}[|X_{i}|^{2\nu_{i}}].$
@@
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\psi((X_{i})_{i>k})\right]=\frac{1}{\prod_{j=1}^{k}\Gamma(\nu_{j})}\int_{(0,\infty)^{k}}\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\psi((X_{i})_{i>k})\right]\prod_{j=1}^{k}s_{j}^{\nu_{j}-1}\mathrm{d}\boldsymbol{s},$
@@
Since $\Sigma^{-1}\preceq\Sigma^{-1}+T_{\boldsymbol{s}}$ in the Loewner order, then $(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}\preceq\Sigma$, and thus $\smash{\boldsymbol{Y}^{(\boldsymbol{s})}}$ is smaller or equal to $\boldsymbol{X}$ in the componentwise convex order, written $\smash{\boldsymbol{Y}^{(\boldsymbol{s})}}\leq_{\mathrm{ccx}}\boldsymbol{X}$, by Lemma 2, so that
@@
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\times\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]\leq\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\right]\times\mathsf{E}\left[\psi((X_{i})_{i>k})\right].$
@@
$\mathsf{E}[f(|\boldsymbol{\alpha}\cdot\boldsymbol{U})g(|T\boldsymbol{\alpha}\cdot\boldsymbol{U})]=\mathsf{E}[f(|T\boldsymbol{\beta}\cdot\boldsymbol{U})g(|\boldsymbol{\beta}\cdot\boldsymbol{U})]=\mathsf{E}[f(|\boldsymbol{e}_{1}\cdot\boldsymbol{U})g(|\boldsymbol{e}_{2}\cdot\boldsymbol{U})].$
@@
The second lemma states that multivariate normal random vectors follow the componentwise convex order according to the Loewner order of their covariance matrices; see, e.g., Shaked & Shanthikumar, (2007, Example 7.A.26).
@@
Let $\Sigma_{1},\Sigma_{2}$ be two $d\times d$ positive definite matrices, and suppose that $\boldsymbol{Y}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma_{1})$ and $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma_{2})$. If $\Sigma_{1}\preceq\Sigma_{2}$, then $\boldsymbol{Y}$ is smaller or equal to $\boldsymbol{X}$ in the componentwise convex order, written $\boldsymbol{Y}\leq_{\mathrm{ccx}}\boldsymbol{X}$, meaning that for every componentwise convex function $\psi:\mathbb{R}^{d}\to\mathbb{R}$, we have
@@
Guolie Lan’s research is supported by the National Natural Science Foundation of China (No. 12171335) and the Guangdong Provincial Natural Science Foundation (No. 2023A1515012170). Frédéric Ouimet’s Research Associate position at McGill University is funded through Christian Genest’s research grants (Grant RGPIN-2024-04088 [Natural Sciences and Engineering Research Council of Canada] and Grant 950-231937 [Canada Research Chairs Program]). Frédéric Ouimet’s postdoctoral fellowship is funded through the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2024-05794 to Anne MacKay). Wei Sun’s research is funded through the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2025-06779).

### v2

@@ Abstract
This paper establishes that the opposite Gaussian product inequality (GPI) of the type proved by Russell & Sun, (2022a) in two dimensions, and partially extended to higher dimensions by Zhou et al. , (2024), continues to hold for an arbitrary mix of positive and negative exponents. A general quantitative lower bound is also obtained conditionally on the GPI conjecture being true.
@@
centered Gaussian random vector , convex order , Gaussian product inequality , Loewner order , multivariate normal distribution , opposite Gaussian product inequality , real linear polarization constant
@@ 1 Introduction
The present paper investigates the general mixed-sign problem, thereby significantly generalizing (3) and (4). Except for a narrow unresolved case discussed in Remark 2, the analysis resolves the question through Theorems 1 and 2.
@@ 2 Results
Next, the goal is to generalize (4) by replacing $|X_{d}|^{2\nu_{d}}$ with a product of positive powers of the absolute Gaussian components indexed in $\mathcal{J}^{\complement}$ or, more generally, a convex function of such components.
@@
Let $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma)$ for some positive definite matrix $\Sigma$, and let $\emptyset\neq\mathcal{J}\subseteq\{1,\ldots,d\}$ be given. Assume that $\nu_{j}\in[0,1/2)$ for all $j\in\mathcal{J}$. Then, for any convex function $\psi:\mathbb{R}^{d-|\mathcal{J}|}\to\mathbb{R}$, we have
@@
Since $(x_{i})_{i\in\mathcal{J}^{\complement}}\mapsto(\sum_{i\in\mathcal{J}^{\complement}}\frac{\nu_{i}}{\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}}|x_{i}|)^{2\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}}$ is convex for $\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}\in[1/2,\infty)$, the following corollary is immediate by combining Theorem 2 with the weighted arithmetic mean-geometric mean inequality.
@@
Let $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma)$ for some positive definite matrix $\Sigma$, and let $\emptyset\neq\mathcal{J}\subseteq\{1,\ldots,d\}$ be given. Assume that $\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}\in[1/2,\infty)$. Then, we have
$\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\times\prod_{i\in\mathcal{J}^{\complement}}|X_{i}|^{2\nu_{i}}\right]\leq\mathsf{E}\left[\prod_{j\in\mathcal{J}}|X_{j}|^{-2\nu_{j}}\right]\times\mathsf{E}\left[\left(\sum_{i\in\mathcal{J}^{\complement}}\frac{\nu_{i}}{\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}}|X_{i}|\right)^{2\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}}\right].$
@@
The case $\sum_{i\in\mathcal{J}^{\complement}}\nu_{i}\in(0,1/2)$ remains open. One possible approach would be to exploit the Bernstein function integral representation, $|x|^{2\nu}=\nu\int_{0}^{\infty}(1-e^{-sx^{2}})s^{-\nu-1}\mathrm{d}s/\Gamma(1-\nu)$, which holds for $x\in\mathbb{R}$ and $\nu\in(0,1/2)$; see, e.g., Schilling et al. , (2012, Eq. (1)). However, we were unable to turn this representation into a successful argument. A treatment of this case therefore lies beyond the scope of our paper.
@@ 3 Proofs
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]$ (6)
$\qquad=\frac{1}{\prod_{j=1}^{k}\Gamma(\nu_{j})}\int_{(0,\infty)^{k}}\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]\prod_{j=1}^{k}s_{j}^{\nu_{j}-1}\mathrm{d}\boldsymbol{s},$
@@
$\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\prod_{i=k+1}^{d}|X_{i}|^{2\nu_{i}}\right]$
$\qquad=\int_{(0,\infty)^{d}}\exp(-\boldsymbol{x}^{\top}T_{\boldsymbol{s}}\boldsymbol{x}/2)\,\frac{\exp(-\boldsymbol{x}^{\top}\Sigma^{-1}\boldsymbol{x}/2)}{(2\pi)^{d/2}|\Sigma|^{1/2}}\prod_{i=k+1}^{d}|x_{i}|^{2\nu_{i}}\mathrm{d}\boldsymbol{x}$
$\qquad=\frac{|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}{|\Sigma|^{1/2}}\int_{(0,\infty)^{d}}\frac{\exp\{-\boldsymbol{x}^{\top}(\Sigma^{-1}+T_{\boldsymbol{s}})\,\boldsymbol{x}/2\}}{(2\pi)^{d/2}|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}\prod_{i=k+1}^{d}|x_{i}|^{2\nu_{i}}\mathrm{d}\boldsymbol{x}$
$\qquad=\frac{|(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}|^{1/2}}{|\Sigma|^{1/2}}\,\mathsf{E}\left[\prod_{i=k+1}^{d}|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right],$
@@
$\mathsf{Var}(Y^{(\boldsymbol{s})}_{i})$ $=\boldsymbol{e}_{i-k}^{\top}\{(\Sigma/A)+B^{\top}A^{-1}(A^{-1}+S)^{-1}A^{-1}B\}\boldsymbol{e}_{i-k}$
$\geq\boldsymbol{e}_{i-k}^{\top}(\Sigma/A)\boldsymbol{e}_{i-k}$
$=\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\mathsf{Var}(X_{i}),$
@@
$\mathsf{E}\left[|Y_{i}^{(\boldsymbol{s})}|^{2\nu_{i}}\right]$$\geq\left\{\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\right\}^{\nu_{i}}\frac{\Gamma(\nu_{i}+1/2)}{\sqrt{\pi}}2^{\nu_{i}}\{\mathsf{Var}(X_{i})\}^{\nu_{i}}$
=$=\left\{\frac{(\Sigma/A)_{i-k,i-k}}{\Sigma_{ii}}\right\}^{\nu_{i}}\mathsf{E}[|X_{i}|^{2\nu_{i}}].$
@@
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\times\psi((X_{i})_{i>k})\right]$
$\qquad=\frac{1}{\prod_{j=1}^{k}\Gamma(\nu_{j})}\int_{(0,\infty)^{k}}\mathsf{E}\left[\exp(-\boldsymbol{X}^{\top}T_{\boldsymbol{s}}\boldsymbol{X}/2)\psi((X_{i})_{i>k})\right]\prod_{j=1}^{k}s_{j}^{\nu_{j}-1}\mathrm{d}\boldsymbol{s},$
@@
Since $\Sigma^{-1}\preceq\Sigma^{-1}+T_{\boldsymbol{s}}$ in the Loewner order, then $(\Sigma^{-1}+T_{\boldsymbol{s}})^{-1}\preceq\Sigma$, and thus $\smash{\boldsymbol{Y}^{(\boldsymbol{s})}}$ is smaller or equal to $\boldsymbol{X}$ in the convex order, written $\smash{\boldsymbol{Y}^{(\boldsymbol{s})}}\leq_{\mathrm{cx}}\boldsymbol{X}$, by Lemma 2, so that
@@
$\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\times\psi((X_{i})_{i>k})\right]\leq\mathsf{E}\left[\prod_{j=1}^{k}|X_{j}|^{-2\nu_{j}}\right]\times\mathsf{E}\left[\psi((X_{i})_{i>k})\right].$
@@
$\mathsf{E}[f(|\boldsymbol{\alpha}\cdot\boldsymbol{U})g(|T\boldsymbol{\alpha}\cdot\boldsymbol{U})]$ $=\mathsf{E}[f(|T\boldsymbol{\beta}\cdot\boldsymbol{U})g(|\boldsymbol{\beta}\cdot\boldsymbol{U})]$
$=\mathsf{E}[f(|\boldsymbol{e}_{1}\cdot\boldsymbol{U})g(|\boldsymbol{e}_{2}\cdot\boldsymbol{U})].$
@@
The second lemma states that multivariate normal random vectors follow the convex order according to the Loewner order of their covariance matrices; see, e.g., Shaked & Shanthikumar, (2007, Example 7.A.13).
@@
Let $\Sigma_{1},\Sigma_{2}$ be two $d\times d$ positive definite matrices, and suppose that $\boldsymbol{Y}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma_{1})$ and $\boldsymbol{X}\sim\mathcal{N}_{d}(\boldsymbol{0}_{d},\Sigma_{2})$. If $\Sigma_{1}\preceq\Sigma_{2}$, then $\boldsymbol{Y}$ is smaller or equal to $\boldsymbol{X}$ in the convex order, written $\boldsymbol{Y}\leq_{\mathrm{cx}}\boldsymbol{X}$, meaning that for every convex function $\psi:\mathbb{R}^{d}\to\mathbb{R}$, we have
@@
Guolie Lan’s research is supported by the National Natural Science Foundation of China (No. 12171335) and the Guangdong Provincial Natural Science Foundation (No. 2023A1515012170). Frédéric Ouimet’s previous Research Associate position at McGill University was funded through Christian Genest’s research grants (Grant RGPIN-2024-04088 [Natural Sciences and Engineering Research Council of Canada] and Grant 950-231937 [Canada Research Chairs Program]). Frédéric Ouimet’s previous postdoctoral fellowship was funded through the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2024-05794 to Anne MacKay). Wei Sun’s research is funded through the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2025-06779).